OpenStudy (anonymous):

Will medal ! Last July a hotel had an average of 280 guests a night. This July the number decreased by 12.5%. How many guests did they have this July ?

OpenStudy (egbeach):

whats 280 divided by 100?

OpenStudy (egbeach):

once you find that you multiple that number by 12.5.

OpenStudy (anonymous):

no this july= 280- 280x0.125 = 280-35= 245

OpenStudy (egbeach):

or you could solve it by 280/100= 2.8 2.8x12.5=35 280-35=245 it depends on how you learn it and which is easier for the person

OpenStudy (anonymous):

Thank you both XD But @egbeach I will give you the medal , because you answered first and explained a bit more :)